OpenStudy (anonymous):

the quantity of x plus 16 all over 3 = 3x

OpenStudy (anonymous):

\[ \frac{x+16}{3} = 3x \\ \frac{x+16}{3} = 3x \;//\cdot 3\\ x+16 = 9x \\ 16 = 8x \\ x = 2 \]

OpenStudy (anonymous):

but don't understand how 9x turned to 8x.

OpenStudy (anonymous):

oh, I see. Well, this is an equation. That means, in order for this to be true both sides actually have the same value. So like, they represent the same number. That means that if i subtract a given number from both sides, it remains equal. So just, subtract x from both sides \[ x + 16 = 9x \\ x + 16 - x = 9x -x \\ 16 = 8x \]